OpenStudy (anonymous):

help with 3 parts please The function H(t) = -16t2 + vt + s shows the height H (t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Part B: What is the maximum height that the projectile will reach? Show your work. (2 points)

OpenStudy (anonymous):

Part C: Another object moves in the air along the path of g(t) = 10 + 63.8t where g(t) is the height, in feet, of the object from the ground at time t seconds. Use a table to find the approximate solution to the equation H(t) = g(t) and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A and estimate using integer values] (4 points) Part D: Do H(t) and g(t) intersect when the projectile is going up or down and how do you know? (2 points)

OpenStudy (anonymous):

@agreene

OpenStudy (agreene):

For part A plug in the values given: \[h(t)=-16t^2+vt+s\\for: v=60;~s=82\\h(t)=-16t^2+60t+82\] use that for the rest of the problem set

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

im solving for t right?

OpenStudy (agreene):

right, when the object hits the ground that will be when y in feet is 0

OpenStudy (anonymous):

okay part b?

OpenStudy (anonymous):

@agreene

OpenStudy (anonymous):

@SCARLET_SPIDER

OpenStudy (anonymous):

@juanpabloJR

OpenStudy (anonymous):

@SolomonZelman I really need help with b,c,d